A consumer has the following utility function for good X and good Y 

𝑈 = 4𝑋 + 8𝑌 

The consumer’s income is KShs 12,000, the price of good X is KShs 300 and the price of good Y is KShs 500. 

Required: 

(i) What type of preferences do the two goods represent? Explain  (ii) Compute the marginal rate of substitution of the two goods  (iii) Calculate the corresponding consumer demand functions  


1
Expert's answer
2021-12-09T09:36:12-0500

i). Straight line preference: The utility function is a linear equation implying that the two goods are perfect sabstitutes which are characterised by straight line prefernce.

ii). "Marginal Rate of Substitution (MRS)=\frac{MUX}{MUY}"

"MUX=4"

"MUY=8"

"MRS=\frac{4}{8}=0.5"

iii). Corresponding demand functions

"MRS=0.5"

"\frac{PX}{PY}=\frac{300}{500}=0.6"

"\frac{PX}{PY}>0.5"

Therefore the consumer will not consume Y and the resulting demand function will be

"Y=\frac{M}{PY}= \frac{12,000}{500}=24"

"Y=24"